Pls tell me the Value of 1/(cosA*sinA) = ?!!!!!!!!!

Grade:11

1 Answers

Darsh patel
6 years ago
See there are two ways to solve this answer..........i). 1\sinAcosA= =(1\sinA)*(1\cosA)=cosecA*SecAii). 1\sinAcosA= =on multiplying by 2 and dividing by 2=(2*1)/2*sinAcosA= 2/sin2A=2Cosec2APlz approve my answer and can ask me again if any query exist... With best regards -Darsh Patel